# algorithmic

### v2

@@ 1 Introduction
Online learning of nonlinear parametric models is of paramount importance in several domains, including model-based adaptive control and real-time estimation of unmeasured variables. Typically, parametric models derived from physics [1] or black-box [2] structures are identified offline on training data, then directly deployed and used without any further updates. On the other hand, further adapting the model online can significantly improve its predictive capabilities [3], especially when the phenomenon we are modeling changes over time. Moreover, online model learning allows for smaller model structures that adapt to varying operating conditions, unlike single, overall models trained offline to cover all conditions.
A vast literature currently exists for online learning [4] and several approaches have been investigated, such as stochastic gradient descent (SGD) [5], follow-the-regularized-leader (FTRL) [6], and online ADMM (alternating direction method of multipliers) [7]. While such approaches provide reasonable learning performance with limited computational effort and can deal with quite general loss functions and regularization terms, they usually learn very slowly, which might be a critical issue, for example, in adaptive control applications.
By treating parameters as constant states, extended Kalman filtering (EKF) has also been proven to be an effective strategy for recursively adapting models from measurements [8, 9, 10, 11, 12, 13] with a faster convergence rate [14]. In particular, [12] proposed a modification of the classical EKF to deal with forgetting factors and an exponential moving-average regularization, improving the performance and flexibility of the online parameter estimation setting, while [11] investigated the use of EKF for the online training of neural network models under general smooth convex losses and smooth regularization functions, with $\ell_{1}$-regularization treated as a special limit case of a smooth regularization. In fact, the main limitation of EKF is that it requires quadratic approximations of nonlinear penalty functions to be able to rephrase the penalty as a squared Euclidean norm of a properly-defined measurement error. This limitation makes EKF not directly suitable for dealing with general non-smooth regularization terms, such as $\ell_{0}$ regularization, group-Lasso penalties, and bound constraints on model parameters, which could instead be beneficial to reduce the complexity of the learned model.
@@
In this paper, we propose a simple and computationally efficient modification of the EKF algorithm by intertwining updates based on online measurements and output prediction errors with updates related to ADMM iterations. This modification allows EKF to deal with a broad class of non-smooth regularization terms for which ADMM is applicable, including $\ell_{0}$/$\ell_{1}$ penalties and bound constraints on model parameters. For linear time-varying models and convex regularization terms, we provide a sublinear regret bound that proves the proper behavior of the resulting online learning strategy. The proposed method is computationally efficient and numerically robust, making it especially appealing for embedded adaptive control applications.
@@ 2 EKF for online model learning
Given a set of input/output data $(z_{k},y_{k})$, $z\in\mathbb{R}^{n_{z}}$, $y\in\mathbb{R}^{n_{y}}$, $k=0,1,\ldots,N-1$, our goal is to recursively estimate a nonlinear parametric model
@@
which describes the (possibly time-varying) relationship between the input and output signals. In (1), $x\in\mathbb{R}^{n_{x}}$ is the parameter vector to be learned, such as the weights of a feedforward neural network mapping $z$ into $y$, or the coefficients of a nonlinear autoregressive model, with $y$ representing the current output and $z$ a vector of past inputs and outputs of a dynamical system. In order to estimate $x$ and capture its possible time-varying nature, we consider the nonlinear dynamical model
@@
where $x_{k}\in\mathbb{R}^{n_{x}}$ is the update of the vector of parameters after collecting $k$ measurements, $h_{k}(\cdot)=h(k,z_{k};\cdot)$, with $h_{k}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}}$ differentiable for all $k$, $y_{k}^{nl}\in\mathbb{R}^{n_{y}}$ is the measurements vector, and $r_{k}\sim\mathcal{N}(0,R_{k}),q_{k}\sim\mathcal{N}(0,Q_{k})$ are the process and measurement noise that we introduce to model, respectively, measurement errors and variations of the model parameters over time, with covariance matrices $R_{k}=R_{k}^{\prime}\succ 0$, $Q_{k}=Q_{k}^{\prime}\succ 0$. By linearizing model (2) around a value $\overline{x}_{k}$ of the parameter vector, i.e., by approximating $h_{k}(x_{k})\approx h_{k}(\overline{x}_{k})+C_{k}(x_{k}-\overline{x}_{k})$, $[C_{k}]_{i:}=\nabla_{x}h_{ki}(\overline{x}_{k})^{\prime}$, $i=1,\ldots,n_{y}$, we obtain the linear time-varying model
@@
with $\overline{x}_{k}=\hat{x}_{k|k-1}$. The first two updates in (4) are usually referred to as the correction step and the last two as the prediction step. Note that (4) is an EKF for model (2), since $C_{k}$ is the Jacobian of the output function at $\hat{x}_{k|k-1}$ and the output prediction error used in the correction step is $e_{k}=y_{k}^{m}-C_{k}=y_{k}^{nl,m}-h_{k}(\hat{x}_{k|k-1})$.
@@
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}$ (6a)
$\hat{x}_{k+1|k},\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k+1},x_{k}}\left\lVert x_{k}-\hat{x}_{k|k}\right\rVert^{2}_{P_{k|k}^{-1}}+\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{Q_{k}^{-1}}$ (6b)
where $\hat{x}_{k+1|k},\hat{x}_{k|k},P_{k+1|k}$, and $P_{k|k}$ are the state estimates and covariance matrices computed as in (4).
@@ 3 EKF under non-smooth regularization
We want to modify the classical iterations (4) by changing the minimization in (6a) with the following
@@
$\nu^{t+1}$ $=$ $\operatorname*{arg\,min}_{\nu}g(\nu)+\frac{\rho}{2}\left\lVert\nu-\hat{x}_{k|k}^{t+1}-w^{l}\right\rVert_{2}^{2}=\text{prox}_{\frac{g}{\rho}}(\hat{x}_{k|k}^{t+1}+w^{t})$ (9b)
@@
for $t=0,\dots,n_{a}-1$, where $\rho>0$ is a hyper-parameter to be calibrated and “prox” is the proximal operator [18]. As shown in [17], in the convex case, the ADMM iterations (9a)–(9c) converge to the optimizer of (8) as $n_{a}\rightarrow\infty$, and often converge to a solution of acceptable accuracy within a few tens of iterations.
Iteration (9c) is straightforward to compute; iteration (9b) can be solved explicitly and efficiently with complexity $\mathcal{O}(n_{x})$ for a wide range of non-smooth and non-convex regularization functions $g$, such as $g(x)=\|x\|_{0}$, $g(x)=\|x\|_{1}$, and the indicator function $g(x)=0$ if $x_{\rm min}\leq x\leq x_{\rm max}$ or +$\infty$+ ∞ otherwise [18]. Iteration (9a) can be rewritten as
@@
3.1 Computational complexity
Given the block-diagonal structure of the measurement noise covariance matrix $\overline{R}_{k}$, in (10), we can separate the contributions of the true measurements $y_{k}^{m}$ and of the fake regularization measurements $\nu^{t}-w^{t}$; moreover, we can process the measurements $\nu^{t}_{i}-w^{t}_{i}$ separately one by one. This allows designing a computationally more efficient and numerically robust version of the proposed EKF-ADMM algorithm, as the correction due to the true measurements $y_{k}^{m}$ can be performed only once, instead of $n_{a}$ times, and there is no need for any matrix inversion when processing the fake measurements. Assuming a complexity $\mathcal{O}(n_{x})$ for evaluating the proximal operator, EKF-ADMM has complexity $\mathcal{O}(n_{x}^{3}+n_{a}n_{x}^{2})$, which is the same order of the full EKF for general state estimation. Moreover, EKF-ADMM has the same number of Jacobian matrices evaluations than the classical EKF, which is usually the most time-consuming part in case $x$ represents the weights and bias terms of a neural network model to learn. Summarizing, the proposed approach is computationally efficient and, if the Kalman filter is implemented using numerically robust factored or square-root modifications [19, 20], the method is appealing for embedded applications.
@@ 4 Regret analysis
     $\nu_{k+1}\leftarrow&\operatorname*{arg\,min}_{\nu}g(\nu)+w_{k}^{T}(x_{k+1}-\nu)+\frac{\rho}{2}\left\lVert x_{k+1}-\nu\right\rVert^{2}=\text{prox}_{\frac{g}{\rho}}(x_{k+1}+\frac{w_{k}}{\rho})\end{aligned}$
@@
The following Theorem 4.1 is an extension of [7, Theorem 4], and provides conditions for sublinear regret bounds of Algorithm 2 in the case of a linear time-varying model (11) and convex regularization function $g$.
@@
Let $\{x_{k},\nu_{k},w_{k}\}_{k=0}^{N-1}$ be the sequence generated by Algorithm 2 and let
$x^{\star},\nu^{\star}=\operatorname*{arg\,min}_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$
be the best solution in hindsight. Let the following assumptions hold:
@@ 5.1 Online LASSO
Consider the LASSO problem $\min_{x}\sum_{k=0}^{N-1}(\frac{1}{2}\left\|y_{k}^{m}\right.$- $\left.C_{k}x\right\|^{2}_{R^{-1}}$ $+$ $\lambda\left\lVert x\right\rVert_{1})$, where $x\in\mathbb{R}^{3}$ is the parameter vector, $C_{k}\in\mathbb{R}^{2\times 3}$ are randomly generated matrices with coefficients drawn from the standard normal distribution, and $y_{k}^{m}=C_{k}x_{\rm true}+r_{k}\in\mathbb{R}^{2}$ is the vector of measurements. We will evaluate the behavior of the regret functions $R_{f}(N)$ and $R_{c}(N)$ as $N\rightarrow\infty$ when using Algorithm 2. The following EKF-ADMM settings are used: $P_{0|-1}=I,Q_{k}=10^{-6}I,R=10^{-3}I,k_{n}=10^{3},\rho=10^{4}\sqrt{N}$ and $\eta=10^{-6}\sqrt{N}$. Results for different values of $\lambda$ and $N$ are shown in Figure 1. In this case, Theorem 2 holds and, as expected, both the regrets $R_{f}(N)$ and $R_{c}(N)$ decrease as the number $N$ of samples increases.
@@ 5.2 Online learning of a static model
Consider the dataset generated by the static nonlinear model
$y^{m}_{k}=\frac{z_{1,k}^{2}-e^{\frac{z_{2,k}}{10}}}{3+|z_{1,k}+z_{2,k}|}+r_{k}.$
We want to train online a neural network $h_{k}(x)$ with $2$ layers, $8$ neurons in each layer, and $\tanh$ activation function, with $n_{x}=105$ trainable weights in total. The training is performed on $N=10^{5}$ randomly generated data points. Let $\{x_{k}\}_{k=0}^{N-1}$ be the sequence of weights generated by Algorithm 1. We evaluate the performance by means of the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-h_{k}(x)\right\rVert^{2}$ and the optimal solution in hindsight is computed by performing 150 epochs using the NAILM algorithm proposed in [23]. The following performance indices will be used for evaluating the quality of the current solution $x$:
$\begin{array}[]{rl}\text{Loss}(x)&=\frac{1}{N}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))\\ \text{Mse}(x)&=\frac{1}{N}\sum_{k=0}^{N-1}f_{k}(x)\\ \text{Reg}(x)&=\frac{1}{N}\sum_{k=0}^{N-1}g(x)\\ \text{Cv}(x)&=\left\lVert x-\Pi_{\mathcal{C}}(x)\right\rVert_{2}^{2}\end{array}$
where $\Pi_{\mathcal{C}}(x)$ is the projection of the point $x$ onto the set $\mathcal{C}$. The training is performed in MATLAB R2022a on an Intel Core i7 12700H CPU with 16 GB of RAM, using the library CasADi [24] to compute the required Jacobian matrices via automatic differentiation. All results are averaged over 20 runs starting from different initial conditions, that were randomly generated using Xavier initialization [25].

### v3

@@ 1 Introduction
Online learning of nonlinear parametric models is of paramount importance in several domains, including model-based adaptive control and real-time estimation of unmeasured variables. Typically, parametric models derived from physics [1] or black-box [2] structures are identified offline on training data, then directly deployed and used without any further updates. On the other hand, further adapting the model online can significantly improve its predictive capabilities [3], especially when the phenomenon we are modeling changes over time, and allows for smaller model structures that adapt to varying operating conditions, unlike single, overall models trained offline to cover all conditions.
A vast literature exists for online learning [4], and several approaches using first or second order gradient information are suitable for real-time model adaptation. Examples of algorithms exploiting first-order information are online gradient descent (OGD) and its regularized alternatives, i.e., follow the regularized leader (FTRL) and online mirror descent (OMD) [5], which are usually fast at execution but slow at convergence. Second-order gradient information can be used to improve the convergence speed, such as in the online Newton step (ONS), or in the well known extended Kalman filter (EKF) [6] which has been proved to be very effective in online model adaptation by treating the parameters as constant states to be estimated [7, 8, 9].
While FTRL and OMD, as well as modified versions of the standard EKF [8], can effectively deal with smooth regularization terms, non-smooth regularizers are often required when learning models. Non-smooth $\ell_{0}$ or $\ell_{1}$ penalties are used to induce sparsity in the model [10], and group-Lasso terms to remove entire parts of the model, such as neurons of a neural network [11]. Indeed, obtaining compact models is particularly important for embedded applications, such as model predictive control, where the real-time numerical complexity of the controller depends directly on the complexity of the prediction model. In addition, indicator functions of feasible sets are another example of non-smooth penalties that are used to impose constraints on model coefficients, such as known bounds on certain unknown physical parameters [12].
Several approaches have been proposed for online learning under non-smooth regularization. Online ADMM (alternating direction method of multipliers) [13] allows handling quite general non-smooth regularizers, while EKF can be modified to deal with $\ell_{1}$-regularization by treating it as a special limit case of a smooth regularizer [8]. While the main limitation of online ADMM is its convergence speed, the main limitation of EKF is that it is not directly suitable for dealing with general non-smooth regularization terms, such as $\ell_{0}$ regularization, group-Lasso penalties, and indicator functions of feasible sets. The main contribution of this work is to show interesting connections and similarities between these two approaches, developing an extension of EKF that inherits the advantages from both, i.e., the convergence speed of EKF and the capability to handle general non-smooth regularization terms of online-ADMM.
@@
The proposed method consists of a simple and computationally efficient modification of the EKF algorithm by intertwining updates based on online measurements and output prediction errors with updates related to ADMM iterations. This modification allows EKF to deal with a broad class of non-smooth regularization terms for which ADMM is applicable, including $\ell_{0}$/$\ell_{1}$ penalties, group-Lasso and indicator functions of simple sets. For linear time-varying models and convex regularization terms, we provide a sublinear regret bound that proves the proper behavior of the resulting online learning strategy. The proposed method is computationally efficient and numerically robust, making it especially appealing for embedded adaptive control applications.
@@
1.1 Notation
Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i:}$ denotes the ith row of A, $A_{:j}$ its $j$th column, $A_{ij}$ its $(i,j)$th entry. Given a vector $v\in\mathbb{R}^{n}$ we denote by $v^{m}$ a measurement of $v$, by $\lVert v\rVert_{1}=\sum_{i=1}^{n}|v_{i}|$ the 1-norm of $v$ and by $\lVert v\rVert_{0}$ its 0-norm, which is defined as the number of non-zero elements in the vector. Given a symmetric positive semidefinite matrix $P=P^{T}\succeq 0$, $P\in\mathbb{R}^{n\times n}$, we denote by $\lVert v\rVert^{2}_{P}$ the quadratic form $v^{T}Pv$. Further, we denote by $\hat{v}_{i|j}$ the estimate of vector $v$ at instant $i$ given all information up to instant $j$, and by $P_{i|j}$ the corresponding covariance matrix. $v\sim\mathcal{N}(\mu,P)$ and $v\sim\mathcal{U}_{[v_{min},v_{max}]}$ denote that v was randomly generated from a normal distribution with mean $\mu$ and covariance $P$ or from a uniform distribution in the interval $[v_{min},v_{max}]$, respectively.
@@ 2 EKF for online model learning
Given a dataset $(z_{k}^{m},y_{k}^{m})$, $z\in\mathbb{R}^{n_{z}}$, $y\in\mathbb{R}^{n_{y}}$, $k=0,1,\ldots,N-1$, our goal is to recursively estimate a nonlinear parametric model
@@
which describes the (possibly time-varying) relationship between the input signal $z_{k}^{m}$ and the output signal $y_{k}^{m}$. In (1), $x\in\mathbb{R}^{n_{x}}$ is the vector of parameters to be learned, such as the weights of a feedforward neural network mapping $z$ into $y$, or the coefficients of a nonlinear autoregressive model, with $y$ representing the current output and $z$ a vector of past inputs and outputs of a dynamical system. In order to estimate $x$ and capture its possible time-varying nature, we consider the nonlinear dynamical model
@@
where $h_{k}(\cdot)=h(k,z_{k};\cdot)$ and we assume $h_{k}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}}$ differentiable for all $k$, $y_{k}^{nl}\in\mathbb{R}^{n_{y}}$, and $r_{k}\sim\mathcal{N}(0,R_{k}),q_{k}\sim\mathcal{N}(0,Q_{k})$ are the measurement and process noise that we introduce to model, respectively, measurement errors and variations of the model parameters over time, with corresponding covariance matrices $R_{k}=R_{k}^{\prime}\succ 0$, $Q_{k}=Q_{k}^{\prime}\succ 0$. By linearizing model (2) around a value $\overline{x}_{k}$ of the parameter vector, i.e., by approximating $h_{k}(x_{k})\approx h_{k}(\overline{x}_{k})+C_{k}(x_{k}-\overline{x}_{k})$, $C_{k,i:}=\nabla_{x}h_{ki}(\overline{x}_{k})^{\prime}$, $i=1,\ldots,n_{y}$, we obtain the linear time-varying model
@@
with $\overline{x}_{k}=\hat{x}_{k|k-1}$. The first two updates in (4) are usually referred to as the correction step and the last two as the prediction step. Note that (4) is an EKF for model (2), since $C_{k}$ is the Jacobian of the output function at $\hat{x}_{k|k-1}$ and the output prediction error used in the correction step is $e_{k}=y_{k}^{m}-C_{k}\hat{x}_{k|k-1}=y_{k}^{nl,m}-h_{k}(\hat{x}_{k|k-1})$.
@@
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}~{}$ (6a)
$\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{k},x_{k+1}}\left\lVert x_{k}-\hat{x}_{k|k}\right\rVert^{2}_{P_{k|k}^{-1}}+\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{Q_{k}^{-1}}~{}$ (6b)
where $\hat{x}_{k|k},\hat{x}_{k+1|k},P_{k|k}$, and $P_{k+1|k}$ are the state estimates and covariance matrices computed as in (4).
@@ 3 EKF under non-smooth regularization
In order to regularize the model, we modify the classical iterations (4) by changing the minimization in (6a) to
@@
$\nu^{t+1}$ $=$ $\operatorname*{arg\,min}_{\nu}g(\nu)+\frac{\rho}{2}\left\lVert\nu-\hat{x}_{k|k}^{t+1}-w^{l}\right\rVert_{2}^{2}$ (9b)
$=$ $\text{prox}_{\frac{g}{\rho}}(\hat{x}_{k|k}^{t+1}+w^{t})$
@@
for $t=0,\dots,n_{a}-1$, where $\rho>0$ is a hyper-parameter to be calibrated and “prox” is the proximal operator [15]. As shown in [12], in the convex case, the ADMM iterations (9a)–(9c) converge to the optimizer of (8) as $n_{a}\rightarrow\infty$, and often converge to a solution of acceptable accuracy within a few tens of iterations. Iteration (9c) is straightforward to compute; iteration (9b) can be solved explicitly and efficiently with complexity $\mathcal{O}(n_{x})$ for a wide range of non-smooth and non-convex regularization functions $g$, such as $g(x)=\|x\|_{0}$, $g(x)=\|x\|_{1}$, and the indicator function $g(x)=0$ if $x_{\rm min}\leq x\leq x_{\rm max}$ or +$\infty$+ ∞ otherwise [15]. Iteration (9a) can be rewritten as
@@
3.1 Computational complexity
Given the block-diagonal structure of the measurement noise covariance matrix $\overline{R}_{k}$, (10) can be rewritten as
$\hat{x}_{k|k}^{t+1}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\sum_{i=1}^{n_{x}}\left\lVert x_{k,i}-\nu^{t}_{i}+w^{t}_{i}\right\rVert_{2}^{2}$ (11)
which highlights the separation of the contributions of the true measurements $y_{k}^{m}$ and of the fake regularization measurements $\nu^{t}-w^{t}$ in the correction step; moreover, we can process the measurements $\nu^{t}_{i}-w^{t}_{i}$ separately one by one. This allows designing a computationally more efficient and numerically robust version of the proposed EKF-ADMM algorithm, as the correction due to the true measurements $y_{k}^{m}$ can be performed only once, instead of $n_{a}$ times, as it does not change with $t$. Moreover, there is no need for any matrix inversion when processing the fake measurements, as by processing them one by one, the matrix inversion required to compute the Kalman gain becomes a simple division, since each measurement is just a scalar value. Assuming a complexity $\mathcal{O}(n_{x})$ for evaluating the proximal operator, EKF-ADMM has complexity $\mathcal{O}(n_{x}^{3}+n_{a}n_{x}^{2})$, which is the same order of the full EKF for general state estimation. Moreover, EKF-ADMM has the same number of Jacobian matrices evaluations than the classical EKF, which is usually the most time-consuming part in case $x$ represents the weights and bias terms of a neural network model to learn. Summarizing, the proposed approach is computationally efficient and, if the Kalman filter is implemented using numerically robust factored or square-root modifications [16], the method is appealing for embedded applications.
@@ 4 Regret analysis
     $\nu_{k+1}\leftarrow&\operatorname*{arg\,min}_{\nu}g(\nu)+w_{k}^{T}(x_{k+1}-\nu)+\\ &+\frac{\rho}{2}\left\lVert x_{k+1}-\nu\right\rVert^{2}=\text{prox}_{\frac{g}{\rho}}(x_{k+1}+\frac{w_{k}}{\rho})\end{aligned}$
@@
The following Theorem 4.1 is an extension of [13, Theorem 4], and provides conditions for sublinear regret bounds of Algorithm 2 in the case of a linear time-varying model (12) and convex regularization function $g$.
@@
Let $\{x_{k},\nu_{k},w_{k}\}_{k=0}^{N-1}$ be the sequence generated by Algorithm 2 and let $x^{\star},\nu^{\star}$ be the best solution in hindsight, i.e. $x^{\star},\nu^{\star}=\operatorname*{arg\,min}_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$. Let the following assumptions hold:
@@ 5.1 Online LASSO
Consider the LASSO problem $\min_{x}\sum_{k=0}^{N-1}(\frac{1}{2}\left\|y_{k}^{m}\right.$- $\left.C_{k}x\right\|^{2}_{R^{-1}}$ $+$ $\lambda\left\lVert x\right\rVert_{1})$, where $x\in\mathbb{R}^{3}$ is the parameter vector, $C_{k}\in\mathbb{R}^{2\times 3}$ are randomly generated matrices with $C_{k,ij}\sim\mathcal{N}(0,1)$, $y_{k}^{m}=C_{k}x_{\rm true}+r_{k}\in\mathbb{R}^{2}$ is the vector of measurements and $r_{k}$ is random measurement noise, $r_{k,i}\sim{\mathcal{N}}(0,10^{-2})$. We will evaluate the behavior of the regret functions $R_{f}(N)$ and $R_{c}(N)$ as $N\rightarrow\infty$ when using Algorithm 2. The following settings are used: $P_{0|-1}=I,Q_{k}=10^{-6}I,R=10^{-3}I,k_{n}=10^{3},\rho=10^{4}\sqrt{N}$ and $\eta=10^{-6}\sqrt{N}$, where $P_{0|-1},Q_{k}$ and $R$ are manually tuned to maximize the Kalman filter performance without regularization and $\rho,\eta$ are chosen according to the expressions provided in Theorem 4.1. Results for different values of $\lambda$ are shown in Figure 1. In this case, Theorem 2 holds and, as expected, both the regrets $R_{f}(N)$ and $R_{c}(N)$ decrease as the number $N$ of samples increases.
@@ 5.2 Online learning of a static model
Consider the dataset generated by the static nonlinear model $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$. We want to train online a neural network $h_{k}(x)$ with $2$ layers, $8$ neurons in each layer, and $\tanh$ activation function, with $n_{x}=105$ trainable weights in total. The training is performed on $N=10^{5}$ randomly generated data points, where $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. Let $\{x_{k}\}_{k=0}^{N-1}$ be the sequence of weights generated by Algorithm 1: we evaluate the online adaptation performance by means of the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-h_{k}(x)\right\rVert^{2}$, and the quality of a given solution $x$ using the performance indices $\text{Loss}(x)=\frac{1}{N}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, $\text{Mse}(x)=\frac{1}{N}\sum_{k=0}^{N-1}f_{k}$, $\text{Reg}(x)=\frac{1}{N}\sum_{k=0}^{N-1}g(x)$ and $\text{Cv}(x)=\left\lVert x-\Pi_{\mathcal{C}}(x)\right\rVert_{2}^{2}$, where, given a constraint set $\mathcal{C}\subseteq\mathbb{R}^{n_{x}}$, $\Pi_{\mathcal{C}}(x)$ is the projection of the point $x$ onto $\mathcal{C}$. The training is performed in MATLAB R2022a on an Intel Core i7 12700H CPU with 16 GB of RAM, using the library CasADi [19] to compute the required Jacobian matrices via automatic differentiation. All results are averaged over 20 runs starting from different initial conditions, that were randomly generated using the well known Xavier’s initialization strategy.